Two cars A and B are running with same speed v .A along east and B towards north .Car A is at origin O and B is at distance d as shown .After how much time  , the distance between cars is minimum and what us the minimum distance?

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Kindly post the image for the position of cars as mentioned

Regards
Khimraj
4 years ago
Let distance covered by A be S1
Let distance covered by B be S2
B will catch A when S1 = S2
S = ut + 1/2 * a * t^2
Therefore, 
40t + (1/2 * 0 * t^2) = 0t + 1/2 * 4 * t^2
40t = 2t^2
20t = t^2
Therefore, t = 20